OpenStudy (anonymous):

salesman earned a 6% commison on sales, and in one week he earned $300 in commissions, how much merchandise did he sell that week

OpenStudy (anonymous):

well okay this is pretty simple want me to guide you through it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so to solve this we first need to know our FACTS Our facts for this problem are: - he earns a 6% commission on sales - he earned $300 in commission what we WANT to know: - how much merchandise did he sell that week This is really simple if you learn how to do it, which you will, don't worry So 6% is sort of awkward if you want to use it in a problem, because of the % sign. Let's convert it to decimal 6% = 6/100 6% = 0.06 Ok that already looks so much better and we are actually almost done! Now all you need to do is remember "he earned $300 in commission". So let's take back the 0.06 Now, to find out how much he sold, let's divide 300 by 0.06 let x = the amount of merchandise he sold 300/0.06 = x 300/0.06 = 5000 x = 5000 So he sold $5000 in merchandise that week :) Want me to post the simpler version?

OpenStudy (anonymous):

Also if you want to check this and make sure that it is true find 6% of $5000 5000 * 0.06 = 300 So yes, this is right!

OpenStudy (anonymous):

@kfishburne26